OpenStudy (anonymous):

1/2+less than or equal signe -2 or t - 4 greater than or equal signe -3

OpenStudy (anonymous):

1/2\[\le-2 or t - 4\ge-3\]

OpenStudy (anonymous):

\[ok sorry the problem reads 1/2 + \le -2 or t - 4 \ge -3\]

OpenStudy (anonymous):

On an "or" statement you just solve each side. I believe you are saying -1/2t <=-2 or t-4 >=-3.... To solve that you solve each equation on its own. So to get rid of -1/2 (because it is by the t) you multiply both sides by (-2/1) which is the reciprocal. So -2 * -2 = 4 so the left side says t >= 4 **** The reason I changed the <= into a >= is because when you multiply or divide both sides by a negative number the rule says you have to flip the signs. The second equation t-4 >= -3 to get t by itself, you add 4 to both sides. So you would have t >= 1 So if your answers are t>= 4 or t >=1 and you graphed those you would find that any number >= 1 would make this true. So your final answer should be t >=1 ** This is a tricky question, if you need more explanation... let me know